The sum of Rex's age and his son, Rodney's age is 50. Five years ago Rex was three times as old as Rodney. What are their ages now?

Answer by Paul(988)

Rex's age = x
Rodney = y

x+y=50
y=50-x (sub)

FIve years ago rex's age = x-5
Five years ago son's age = y-5

3(y-5)=x-5
3y-5=x-5
SUb
3(50-x-5)=x-5
135-3x=x-5
-4x=-140
x=35

y=50-35
y=15

Hence, Rex is 35 years old and his son is 15 years old.
PAul.